OpenStudy (anonymous):

someone help me? another medal and fan if we can work to find out the answer!

OpenStudy (anonymous):

Calculate the amount of your compound interest investment after 10 years. Remember that you are starting with $7,500. (my interest rate is .84%)

OpenStudy (anonymous):

@satellite73

OpenStudy (psymon):

How many compounds per year? Or is it continuously?

OpenStudy (anonymous):

use the compound interest formula and plug in your numbers.

OpenStudy (anonymous):

the frequency is daily

OpenStudy (psymon):

\[A = P(1+\frac{ r }{ n })^{nt} \] where P is starting amount, r is the rate as a decimal, n is the amount of compounds per year and t is time in years.

OpenStudy (psymon):

Okay, so that means n in our formula would be 365. So then we just plug in numbers: \[A = 7500(1+\frac{ .0084 }{ 365 })^{365*10} \] Assuming you meant .84% and not 84% xD

OpenStudy (anonymous):

i bet n =12

OpenStudy (psymon):

Even though its daily?

OpenStudy (anonymous):

then your daily interest is (0.84%)/365 = 0.84/(100*365) 7500*(1+0.84/(100*365))^(10*365)

OpenStudy (anonymous):

Oh daily sorry i missed that. You are correct Simon

OpenStudy (psymon):

What zarax put and what I put are the same thing, btw @xoxo_devvie, he just had an extra division by 100 in the bottom and I moved the decimal place over 2 to the left without throwing in the extra divide by 100 in the bottom.

OpenStudy (anonymous):

i am lost didn't we just do this exact problem?

OpenStudy (anonymous):

i forgot to tell you my interest rate.

OpenStudy (anonymous):

can you help me work on the rest of the problem @Psymon ?

OpenStudy (anonymous):

@zarax can you hee igure out the rest of the problem?

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

oh hi there c:

OpenStudy (anonymous):

hi! ahahah

zepdrix (zepdrix):

\[\Large A=P(1+\frac{r}{n})^{nt}\] And we have:\[\Large r=.84\text{%} \qquad\to\qquad r=.0084\]\[\Large t=10\]\[\Large n=365\]\[\Large P=7500\]

zepdrix (zepdrix):

Just plug em in!! :O

OpenStudy (anonymous):

ok hold on dont go anywhere! ahah

OpenStudy (anonymous):

my calculter says error

zepdrix (zepdrix):

You silly billy -_-

OpenStudy (anonymous):

ahhh! ahah i dont know what im doing wrong:(

zepdrix (zepdrix):

If you're trying to put it all into your calculator at the same time, try doing it like this:\[\Large 7500((1+(.0084/365))\text{^}(365*10)\]

zepdrix (zepdrix):

I find it really difficult to do it all at once, since you need so many brackets. I prefer to do it piece by piece. But that requires a little bit of finesse :)

OpenStudy (anonymous):

ok hold on haha!

OpenStudy (anonymous):

uuuuughhh! im getting so mad! i got 27375630

zepdrix (zepdrix):

Did you try entering it the way I posted? D: You should get something a lot nicer than that. What kind of calculator are you using?

OpenStudy (anonymous):

just a basic calculater and yes i did!:(

zepdrix (zepdrix):

Do you have a key called \(\large Ans\) anywhere on it?

OpenStudy (anonymous):

nooooo

zepdrix (zepdrix):

Ok then I guess the way we would have to do this is by calculating it piece by piece and rounding between each step :)

zepdrix (zepdrix):

Start with this :)\[\Large .0084\div 365=?\]

OpenStudy (anonymous):

2,301?

zepdrix (zepdrix):

whut? +_+

zepdrix (zepdrix):

You should get \(\Large 0.000023014\), yes?

OpenStudy (anonymous):

whattt? let me try again

OpenStudy (anonymous):

nope i didnt get that

OpenStudy (anonymous):

2.301369 is what i got

zepdrix (zepdrix):

Oh lol. Did you get 2.301369 and then on the far right it has something about an E?

zepdrix (zepdrix):

Something like this maybe? :o\[\large 2.301369 \qquad -5\text{E}\]

OpenStudy (anonymous):

yes!

zepdrix (zepdrix):

That sucks :( your calculator is putting it into `scientific notation` for you. That's going to make this a little more difficult. Maybe use Google Calculator or something for now.

zepdrix (zepdrix):

https://www.google.com/#q=.0084%2F365

OpenStudy (anonymous):

ok so what do i need to do?

zepdrix (zepdrix):

Let's just type the calculations into google search :) .0084/365 like that.

zepdrix (zepdrix):

So we get this long decimal. \(\Large 0.00002301369\) Let's cut it off at some convenient point so it's easier to use.\[\Large 0.000023\]

OpenStudy (anonymous):

ok now thats what i got also!

zepdrix (zepdrix):

Ok good c: So we started with this, \[\Large A=7500(1+\frac{.0084}{365})^{365*10}\]And making our first calculation gives us this,\[\Large A=7500(1+0.000023)^{365*10}\]

zepdrix (zepdrix):

Let's work from the inside out, following our order of operations. So let's add the 1 and the long decimal next.

zepdrix (zepdrix):

1+0.000023 = ?

OpenStudy (anonymous):

1.000023 ?

zepdrix (zepdrix):

Ok good!\[\Large A=7500(1.000023)^{365*10}\] Next let's simplify the exponent by doing the multiplication.\[\Large 365*10=?\]

OpenStudy (anonymous):

3650?

zepdrix (zepdrix):

K cool :)\[\Large A=7500(1.000023)^{3650}\] See how it's getting a little bit simpler as we do each calculation? This problem would be A LOT easier if you had a decent calculator ^^ lol

OpenStudy (anonymous):

oh gosh i know right but thank you for helping me so far!

zepdrix (zepdrix):

For the next calculation we'll use the caret symbol for the exponent operation. 1.000023^3650

OpenStudy (anonymous):

wheres that sign at?

zepdrix (zepdrix):

Shift 6

OpenStudy (anonymous):

1.08757346377 ?

zepdrix (zepdrix):

Ok that sounds right.\[\Large 1.08757346377\] Let's cut off some of the numbers so it's easier to work with. (We'll lose a little bit of accuracy by doing this, but whatever.)\[\Large 1.08757\]

zepdrix (zepdrix):

So now our problem becomes,\[\Large A=7500*1.08757\]Right?

OpenStudy (anonymous):

right!!!

zepdrix (zepdrix):

So what do you get when you do the final multiplication? :o

OpenStudy (anonymous):

8156.775 ?

zepdrix (zepdrix):

If we had punched all the numbers in at the same time we would have gotten,\[\Large A=8157.2088\] See how our answer is off by a tiny bit? It's not a big deal. But that's just something to keep in mind. When we do rounding in the middle of the problem, it will affect our final answer a little bit.

OpenStudy (anonymous):

thats ok! well what would i have gotten?

zepdrix (zepdrix):

what? +_+

OpenStudy (anonymous):

for the actual answer?

zepdrix (zepdrix):

The answer you came up with is fine c: It's close enough to be correct. You could round it up to 8157 if you wanted to.

zepdrix (zepdrix):

Spend 10 bucks, get a decent calculator!! D: It'll make these problems a lot easier :c

zepdrix (zepdrix):

Unless you're not allowed to use calculators on tests..

OpenStudy (anonymous):

its not a test lol its a question on my homework :p would you mind helping me with 2 more problems?:) ill ask it in a whole different question so i can give you medals for those too. you have no idea how happy i am! thank you thank you thank you!

zepdrix (zepdrix):

Ya I'll try to stop by :d

OpenStudy (anonymous):

ok! yay!